OpenStudy (anonymous):

Simplify 3x + (-12x) - 5x. 10x?

hartnn (hartnn):

how did u come up with 10 ?

OpenStudy (anonymous):

Added them together.

hartnn (hartnn):

so, lets do that 3+ (-12)+ (-5) = 3-12-5 = .... ?

OpenStudy (anonymous):

-14x?

hartnn (hartnn):

yes!